1. Current Ratio - Current assets/Current liabilities
Answer #1

Starbucks corporation:

particular 2017 2016
Total Current Assets 5283.40 4757.90
Total Current Liabilites 4220.70 4546.80

Current Ratio: Current Asset/Current Liabilities

2016:- 4757.90/4546.80= 1.046

2017:- 5283.40/4220.70= 1.251

The Classic Cofee Corporation

particular 2017 2016
Total Current Assets 1773.63 1288.54
Total Current Liabilites 611.89 554.20

Current Ratio: Current Asset/Current Liabilities

2016: 1288.54/554.20=2.32

2017: 1773.63/611.89=2.89